OpenStudy (anonymous):

express 3 log7 p - 2 log7 q as a single logarithm

OpenStudy (cwrw238):

3log7 p = log7 p^3 2log7 q = log7 q^2 now use the law log a - log b = log (a/b)

OpenStudy (anonymous):

log7p^3 - log7q^2 = log7p^3 / log7q^2 ?

OpenStudy (cwrw238):

no log7 (p^3/q^2)

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

thanks